Rewrite the expression 16/8log(x+5)+16/4log(x^2+3)-4/8log(x+2) as a single logarithm


1
Expert's answer
2022-06-13T15:19:32-0400
"\dfrac{16}{8}\log(x+5)+\dfrac{16}{4}\log(x^2+3)-\dfrac{4}{8}\log(x+2)"

"=2\log(x+5)+4\log(x^2+3)-\dfrac{1}{2}\log(x+2)"

"=\log\dfrac{(x+5)^2(x^2+3)^4}{\sqrt{x+2}}"